# GHOST commissioning science results - IV: Chemodynamical analyses of Milky Way satellites Sagittarius II and Aquarius II

### v1

@@ 1 Introduction
With the advent of extensive and deep photometric surveys111Photometric surveys such as the Sloan Digital Sky Survey (SDSS: York, 2000; Abazajian et al., 2009), the Dark Energy Survey (DES: Abbott et al., 2018), DELVE (Drlica-Wagner et al., 2021), Panoramic Survey Telescope and Rapid Response System (PS1: Chambers et al., 2016), UNIONS (see Jensen et al., 2021; Smith et al., 2023), and Euclid (e.g., Hunt et al., 2024), it has become possible to discover extremely faint satellites of the Milky Way with . Identifying the physical nature of these new systems has proven to be a challenge, (e.g., UNIONS1/UMaIII, Smith et al. 2023; Cerny et al. 2023; Eridanus III and DELVE 1, Simon et al. 2024)
Some of the new systems are likely ultra faint dwarf galaxies (UFDs), with (Simon, 2019). UFDs provide valuable insights into the faint end of the galaxy luminosity function, galaxy formation models, ancient star formation histories, and chemical evolution in the early universe (Koposov et al., 2009; Starkenburg et al., 2013; Frebel & Norris, 2015). The abundance and distribution of dark matter in the halos of these dim, dense systems are also vital for constraining and the nature and properties of dark matter and cosmological models (Springel et al., 2008; Brooks & Zolotov, 2014; Bullock & Boylan-Kolchin, 2017; Sales et al., 2022). Alternatively, some of the new systems are likely to be (potentially disrupted) star clusters (e.g., Malhan & Ibata, 2018; Ji et al., 2020a; Li et al., 2022a; Martin et al., 2022), Faint star clusters are also valuable probes but for different scientific purposes, e.g., as tracers of their host galaxy properties, ranging from the host’s gravitational potential and assembly history, to its star formation and chemical enrichment history (Kravtsov & Gnedin, 2005; Mészáros et al., 2015; Gaia Collaboration et al., 2018; Li et al., 2019)
@@
Using high-resolution spectra from the newly commissioning Gemini/GHOST spectrograph McConnachie et al. (2024); Kalari et al. (2024), we revisit the analyses of Sgr2 and Aqu2. Our targets include two previously confirmed members of Sgr2, one new member of Sgr2, and two of the brightest stars in Aqu2. Our objectives include: (i) refining constraints on the velocity and metallicity dispersions of these systems, (ii) performing a detailed chemical analysis for the first time in these two faint systems, and (iii) increasing the observational epochs per star to assess any binary characteristics. These endeavors are to contribute to the discussion on the classification of these ambiguous systems as dark matter dominated UFDs or faint stellar clusters.
@@ 2.3 Radial Velocities ()
Radial velocities were determined using cross-correlation (IRAF/fxcor) with the GHOST spectrum of the standard star HD 122563, commonly used as a benchmark for metal-poor stars. The spectral region from 3800 to 6700 Å was used, as the SNR worsens at shorter wavelengths and the sky/telluric lines are poorly correlated at longer wavelengths. Our radial velocities per target are given in Table 2. The for our GHOST spectrum of HD122563 is km s-1 (see Hayes et al., 2023), in excellent agreement with Gaia DR3 and results in the literature. All target are measured from template fitting using this standard star as a reference.
@@ 3 Method
The spectrral analysis was conducted using a new set of jupyter notebooks, Py_Looper444https://github.com/dashazaremba/PyLooper, which offer a semi-automated routine for high-resolution spectral analysis based on equivalent widths (EW). This process starts by determining the metallicity ([Fe/H]) from iron lines, using an initial set of stellar parameters based on photometric calibrations (see Section 3.2). Spectral parameters () can be refined adopting both 1DLTE and 1DNLTE analyses. Subsequently, the routine measures other spectral lines, applying necessary corrections, including NLTE corrections from precalculated grids (see Section 3.5.2), and hyperfine structure corrections555HFS corrections were calculated for odd-Z elements only. For elements where isotopic ratios depend on r/s-process dominance (e.g., Ba II, Eu II), HFS corrections were calculated directly from spectrum syntheses. (HFS) for odd-Z elements. Overall, this method includes error propagation from uncertainties in metallicity and the derived stellar parameters throughout the analysis. For the line measurement errors, we use the line-to-line scatter in Fe I for the other elements as this is a good indicator of impact of the SNR. To validate the Py_Looper method, we conducted a line-by-line comparison of EWs and abundances for Fe lines in the HD222925 standard star (see the Appendix) against the literature values reported by Roederer et al. (2018).
@@
To reduce systematic errors, we compare results for our target stars to two standard stars with well-established stellar parameters, HD122563 and HD222925, which differ in metallicity and Teff. The initial parameters were sourced from the literature (Roederer et al., 2018; Giribaldi et al., 2023) and subsequently analyzed using the same methodology applied to our targets for consistency. The final stellar parameters with their associated uncertainties for the standards are presented in Table 2. The same Fe line list was used for both the standards and the targets. For other elements, the lines were first measured for the two standards, with only moderately strong and strong lines retained, as weaker lines could be too contaminated by noise in our spectra. This refined line list was then consistently applied to the target stars.
@@ 3.2 Stellar Parameters
HD122563 | … | … | 121212Hayes et al. (2023) | |||||
HD222925 | … | … |
@@ 3.3 Spectral Lines Analysis
Our spectral analysis is based on the line list provided in the Appendix. This line list was compiled from our analyses of two standard stars, HD122563 (Venn et al., 2025) and HD222925 (Hayes et al., 2023). Spectral lines from a variety of sources (Yong et al., 2013; Norris et al., 2017; Kielty et al., 2021; Lucchesi et al., 2022; Roederer et al., 2022; Sestito et al., 2024) were examined in these standard stars and their equivalent widths (EWs) compared to those in the literature. All atomic data was adopted from the linemake compilation131313Linemake available at https://github.com/vmplacco/linemake (Placco et al., 2021).
@@ 3.5 Other Elements
Spectral lines of C to Eu are available for chemical abundance measurements in this sample. Some elements requires spectrum syntheses or additional corrections; e.g., due to isotopic splitting, hyperfine structure corrections, and/or NLTE corrections. We do not consider 3D effects in this paper.
@@ 3.5.2 NLTE corrections
NLTE corrections for Ca, Mg, Ti, Mn, Si, and Fe are obtained from the MPIA webtool database141414http://nlte.mpia.de. Corrections for Na are from the INSPECT database151515http://inspect-stars.com (Lind et al., 2011). A python wrapper to extract these corrections for our stars is available161616https://github.com/anyadovgal/NLTE-correction. For K, we apply the NLTE correction grid from Reggiani et al. (2019). Ba NLTE corrections are from Mashonkina & Belyaev (2019). All NLTE line corrections are presented in Table 7 (Fe lines) and Table 8 (other elements) in the Appendix.
@@ 4 Chemical Abundances
CH171717C abundances from CH after applying an evolutionary correction from Placco et al. (2014) | -0.16 0.30 | -0.32 0.30 | -0.22 0.30 | -0.21 0.10 | 0.00 0.50 |
@@ 5.2 Chemistry in Aqu2
We compare our chemical abundances for Aqu2472 to Population III model yields from Heger & Woosley (2010, 2012 update). In Fig. 12, an excellent fit was found when combining the yields from two Pop III SN using StarFit181818StarFit: https://starfit.org/. These two stars included an 18.4 M⊙ SN with typical energy (2.4 x 1051 erg), combined with a 90 M⊙ SN with excess energy (10 x 1051 erg), both with mild mixing fractions. Ti was excluded from the fit in Fig. 12, as its NLTE corrections may be insufficient (e.g., see Mallinson et al., 2022, 2024); including Ti increases the value to 0.82. We note that our Ti abundance can be reproduced by including a third source, particularly with a significant r-process yield; however, the required additional model would also lead to an overproduction of Sr and Ba relative to our upper limit estimates, making such a scenario unlikely. For Aqu2776, none of the tested combinations of 2-3 Pop III SN yields provided a satisfactory fit (with ). Given its proximity to the system’s center, as shown in Fig. 8, and its overall higher metallicity, we attribute this to enrichment from more Pop III stars, and most likely pollution from super-AGB stars (discussed below in Section 5.2.1).
@@ 5.2.1 Element (anti-) correlations (K-Na)
Alternatively, the unusual globular cluste NGC 2419 is notable for its substantial scatter in K abundances, reaching unprecedented values of [K/Fe] dex (Cohen & Kirby, 2012; Mucciarelli et al., 2012); see Fig. 14. These high-K stars are found to be correlated with depleted Mg; i.e., a Mg-K anti-correlation. Ventura et al. (2012) proposed that hot-bottom burning in stars of masses around 6 M⊙ – at the edge between AGB and super-AGB (SAGB) regime – could reproduce the extreme potassium abundances in NGC 2419 if the standard cross section of Argon nuclei in reaction were increased by a factor of 100. In this model191919The SAGB model by Ventura et al. (2012) also predicts a modest increase in Na and a strong depletion in O, i.e., a weak Na-O anti-correlation similar to that observed in GCs. The SAGB scenario has been further supported by recent high-precision measurements of the potassium-destroying reaction (Fox et al., 2024), which is critically sensitive to temperature-density conditions (Dermigny & Iliadis, 2017)., the reduction of Mg and the production of K are maximized by also decreasing the mass loss rate by a factor of 4. By exploring a nuclear reaction network sensitive to variations in temperature, density, hydrogen abundance, reaction rates, and initial composition, Iliadis et al. (2016) identified that these abundance variations are limited to a narrow temperature-density range, achievable only by SAGB stars and classical novae.
@@ 6.2 Chemistry in Sgr2
We also examine the absolute abundances of Sgr2584 and Sgr2936202020Sgr2656 is not plotted due to having fewer abundance measurements, all of which are consistent with the other two targets. compared to solar values and our two standard stars, HD222925 and HD122563. In Fig. 19, the chemistry for the two standard stars in determined in two ways: (1) from the literature for HD222925 (Roederer et al., 2018) and HD122563 (Honda et al., 2006; Collet et al., 2018), and (2) derived using the same spectral lines as for the Sgr2 stars (details discussed further in Venn et al. 2025). The solar system’s s- and r-process abundance patterns from Simmerer et al. (2004) are scaled to match the Ba and Eu abundances in Sgr2584. The chemical abundances for both Sgr2 targets are similar, and generally fall between those of the two standard stars.
@@ 6.3 R-process Enrichments in Stellar Populations
As a reference globular cluster for comparison, M92 exhibits the highest [Eu/Fe] abundances in members located roughly two half-light radii to the northeast; see Fig. 21. However, no clear gradient is observed with respect to radial distance or along the east-west and north-south directions.
@@ 6.4 Comparison of Sgr2 with Tuc III
Tuc III is a particularly interesting system as it has been classified as a UFD based on its low average metallicity ([Fe/H] ; Li et al. 2018), a velocity and metallicity dispersion212121The observed metallicity gradient from CaT lines among core and tidal tail members has likewise been presented as evidence for the system’s origin as a UFD (Li et al., 2018). We reanalyze the evidence for a metallicity dispersion and metallicity gradients in the Appendix and do not find clear evidence for Tuc III to be classified as an UFD., and the absence of light-element anti-correlations amongst five of its brightest members (Hansen et al., 2017; Marshall et al., 2019). However, Marshall et al. (2019) note that adopting the photometric temperature for the metallicity outlier star from Hansen et al. (2017) would eliminate the statistically significant metallicity spread among its members. And, unfortunately, the velocity distribution for Tuc III is not representative of the system due to its disrupted state (only an upper limit of 1.5 km s-1 at the 95.5 % confidence level for core members was derived by Simon et al. 2017). Taking into account that low metallicity is also insufficient evidence for the presence of dark matter (i.e., an UFD classification), as globular clusters are typically found with [Fe/H] , and even lower ([Fe/H] have been found for M31 clusters Simpson 2018; Larsen et al. 2020, and the stellar stream C-19 appears to be a disrupted star cluster with [Fe/H] , see Martin et al. 2022; Yuan et al. 2022; Venn et al. 2025), then Tuc III remains a challenging system to classify.
@@ 7 Additional diagnostics for ambiguous systems
In addition to theoretical uncertainties, the observed velocity values can also be influenced by data quality and sample selection (e.g., compare the pink profile from L20 and the green profile from L21 in Fig. 17), as well as binary contamination. GC observations show that the binary fraction increases toward a cluster center (e.g., Sollima et al., 2007; Ji et al., 2015), and simulations reveal that as binaries segregate, "hard"222222”hard” means that their binding energy is much higher than the average kinetic energy of stars in a cluster binaries become even harder. Their semimajor axes shrink according to the Heggie 1975 - Hills 1975 law, making them more difficult to identify and potentially causing up to a 70% overestimation of the velocity dispersion in the cluster’s core (Aros et al., 2021). As a quantitative example, Wang et al. (2024) estimated the observed binary fraction for a Palomar 5-like globular cluster (which we note is close to Sgr2 on the vs. diagram; see Fig. 1), with a heliocentric distance of 20 kpc. They showed that most bright binaries with periods below days can be detected within 6 months by measuring line-of-sight velocities of km s-1. The undetected binaries ( km s-1) could still inflate the computed velocity dispersion by a factor of 1.5-2 compared to estimates based solely on single stars. For Sgr2, which is times farther away, the stars have measured radial velocity uncertainties 1 km s-1 (L21), constrained by observations of only 1 month (38 days). This implies its () is not resolved. This agrees with Baumgardt et al. (2022), who estimated a lower velocity dispersion for Sgr2232323The velocity dispersion profile for Sgr2 derived with N-body models of Baumgardt (2017) can be found at https://people.smp.uq.edu.au/HolgerBaumgardt/globular/fits/kin/sgrii_vel.pdf of km s-1, i.e., 3x lower than L21 (see Fig. 16), and attributed the higher observed () entirely to binary contamination.
To take these studies further, Baumgardt et al. (2022) have conducted a homogeneous and extensive study of mass segregation in GCs and UFD candidates using N-body simulations. As a measure of mass segregation, they used the ratio of the radii where the cumulative fraction of bright and faint stars reaches 0.5. A lower ratio (deviating further from unity) indicates stronger segregation and is expected in clusters with relaxation times much shorter than their age. To get the two subsets, members below the main sequence turn-off and above a certain MLow mass threshold were divided into two equally sized groups based on brightness. For Sgr2, their mass segregation parameter , which is in agreement with the expected value for a star cluster of Sgr2’s relaxation time of 8.37 Gyr, calculated following Spitzer (1987)242424We note that Wang et al. (2024) find that the presence of black holes can significantly shorten the two-component relaxation time..
@@
We support that the mass segregation parameter can provide an additional and valuable test for low surface brightness systems when trying to determine their origins (GC vs UFD), total mass, and thereby dark matter contents. Such diagnostics are especially important when alternative approaches, like studying system stability within the Milky Way’s tidal field, are inapplicable or computationally expensive (e.g., Sgr2’s distant pericenter, as noted in Section 1, or the impact of the LMC on Tuc III).

### v3

@@ 1 Introduction
With the advent of extensive and deep photometric surveys111Photometric surveys such as the Sloan Digital Sky Survey (SDSS: York, 2000; Abazajian et al., 2009), the Dark Energy Survey (DES: Abbott et al., 2018), DELVE (Drlica-Wagner et al., 2021), Panoramic Survey Telescope and Rapid Response System (PS1: Chambers et al., 2016), UNIONS (see Jensen et al., 2021; Smith et al., 2023), and Euclid (e.g., Hunt et al., 2024), it has become possible to discover extremely faint satellites of the Milky Way with . Identifying the physical nature of these new systems has proven to be a challenge, (e.g., UNIONS1/UMaIII, Smith2024; Eridanus III and DELVE 1, Simon et al. 2024)
Some of the new systems are likely ultra faint dwarf galaxies (UFDs), with (Simon, 2019). UFDs provide valuable insights into the faint end of the galaxy luminosity function, galaxy formation models, ancient star formation histories, and chemical evolution in the early universe (Koposov et al., 2009; Starkenburg et al., 2013; Frebel & Norris, 2015). The abundance and distribution of dark matter in the halos of these dim, dense systems are also vital for constraining the nature and properties of dark matter and cosmological models (Springel et al., 2008; Brooks & Zolotov, 2014; Bullock & Boylan-Kolchin, 2017; Sales et al., 2022). Alternatively, some of the new systems are likely to be (potentially disrupted) star clusters (e.g., Malhan & Ibata, 2018; Ji et al., 2020a; Li et al., 2022a; Martin et al., 2022), Faint star clusters are also valuable probes but for different scientific purposes, e.g., as tracers of their host galaxy properties, ranging from the host’s gravitational potential and assembly history, to its star formation and chemical enrichment history (Kravtsov & Gnedin, 2005; Mészáros et al., 2015; Gaia Collaboration et al., 2018; Li et al., 2019)
@@
Using high-resolution spectra from the newly commissioning Gemini/GHOST spectrograph (McConnachie et al., 2024; Kalari et al., 2024), we revisit the analyses of Sgr2 and Aqu2. Our targets include two previously confirmed members of Sgr2, one new member of Sgr2, and two of the brightest stars in Aqu2. Our objectives include: (i) refining constraints on the velocity and metallicity dispersions of these systems, (ii) performing a detailed chemical analysis for the first time in these two faint systems, and (iii) increasing the observational epochs per star to assess any binary characteristics. These endeavors are to contribute to the discussion on the classification of these ambiguous systems as dark matter dominated UFDs or faint stellar clusters.
@@ 2.3 Radial Velocities ()
Radial velocities were determined using cross-correlation (IRAF/fxcor) with the GHOST spectrum of the standard star HD 122563, commonly used as a benchmark for metal-poor stars. The spectral region from 3800 to 6700 Å was used, as the SNR worsens at shorter wavelengths and the sky/telluric lines are poorly correlated at longer wavelengths. Our radial velocities per target are given in Table 2. The for our GHOST spectrum of HD122563 is km s-1 (see Hayes et al., 2023), in excellent agreement with Gaia DR3 and results in the literature. All target are measured from template fitting using this standard star as a reference. We did not apply any zero-point corrections to the velocity measurements, as our derived values show excellent agreement with those obtained from other spectrographs for the same stars. GHOST’s velocity precision – reaching meters per second (McConnachie et al., 2024; Kalari et al., 2024) – also surpasses that of comparable instruments.
@@ 3 Method
The spectral analysis was conducted using a new set of jupyter notebooks, Py_Looper444https://github.com/dashazaremba/PyLooper, which offer a semi-automated routine for high-resolution spectral analysis based on equivalent widths (EW). This process starts by determining the metallicity ([Fe/H]) from iron lines, using an initial set of stellar parameters based on photometric calibrations (see Section 3.2). Spectral parameters () can be refined adopting both 1DLTE and 1DNLTE analyses. Subsequently, the routine measures other spectral lines, applying necessary corrections, including NLTE corrections from precalculated grids (see Section 3.5.2), and hyperfine structure corrections555HFS corrections were calculated for odd-Z elements only. For elements where isotopic ratios depend on r/s-process dominance (e.g., Ba II, Eu II), HFS corrections were calculated directly from spectrum syntheses. (HFS) for odd-Z elements. Overall, this method includes error propagation from uncertainties in metallicity and the derived stellar parameters throughout the analysis. For the line measurement errors, we use the line-to-line scatter in Fe I for the other elements as this is a good indicator of impact of the SNR. To validate the Py_Looper method, we conducted a line-by-line comparison of EWs and abundances for Fe lines in the HD222925 standard star (see the Appendix) against the literature values reported by Roederer et al. (2018).
@@
To reduce systematic errors, we compare results for our target stars to two standard stars with well-established stellar parameters, HD122563 and HD222925, which differ in metallicity and Teff. The initial parameters were sourced from the literature (Roederer et al., 2018; Giribaldi et al., 2023) and subsequently analyzed using the same methodology applied to our targets for consistency. These parameters result in an excellent fit for our analysis, i.e., no correlations in the FeI line abundances with excitation potential (slope ) or line equivalent widths (slope ), and with good ionization equilibrium between A(FeI) and A(FeII), as described in Section 3.4; thus, we retain these parameters, but adopt our own error analysis. The stellar parameters with their associated uncertainties for the standards are presented in Table 2. The same Fe line list was used for both the standards and the targets. For other elements, the lines were first measured for the two standards, with only moderately strong and strong lines retained, as weaker lines could be too contaminated by noise in our spectra. This refined line list was then consistently applied to the target stars.
@@ 3.2 Stellar Parameters
| HD122563 | … | … | 121212Stellar parameters are from Roederer et al. (2014) | 131313Hayes et al. (2023) | ||||
| HD222925 | … | … | 141414Stellar parameters are from Giribaldi et al. (2023) |
@@ 3.3 Spectral Lines Analysis
Our spectral analysis is based on the line list provided in the Appendix. This line list was compiled from our analyses of two standard stars, HD122563 (Venn et al., 2025) and HD222925 (Hayes et al., 2023). Spectral lines from a variety of sources (Yong et al., 2013; Norris et al., 2017; Kielty et al., 2021; Lucchesi et al., 2022; Roederer et al., 2022; Sestito et al., 2024) were examined in these standard stars and their equivalent widths (EWs) compared to those in the literature. All atomic data was adopted from the linemake compilation151515Linemake available at https://github.com/vmplacco/linemake (Placco et al., 2021).
@@ 3.5 Other Elements
Spectral lines of C to Eu are available for chemical abundance measurements in this sample. Some elements requires spectrum syntheses or additional corrections; e.g., due to isotopic splitting, hyperfine structure corrections, and/or NLTE corrections. We do not consider 3D effects in this paper161616Fully consistent 3D NLTE model atmospheres with line-by-line radiative transfer are not yet available. Furthermore, mean 3D (¡3D¿) NLTE has yet to demonstrate clear advantages over 1D NLTE in individual cases (e.g., Mallinson et al., 2024).
@@ 3.5.2 NLTE corrections
NLTE corrections for Ca, Mg, Ti, Mn, Si, and Fe are obtained from the MPIA webtool database171717http://nlte.mpia.de. Corrections for Na are from the INSPECT database181818http://inspect-stars.com (Lind et al., 2011). A python wrapper to extract these corrections for our stars is available191919https://github.com/anyadovgal/NLTE-correction. For K, we apply the NLTE correction grid from Reggiani et al. (2019). Ba NLTE corrections are from Mashonkina & Belyaev (2019). All NLTE line corrections are presented in Table 7 (Fe lines) and Table 8 (other elements) in the Appendix.
@@ 4 Chemical Abundances
| CH202020C abundances from CH after applying an evolutionary correction from Placco et al. (2014) | -0.16 0.30 | -0.32 0.30 | -0.22 0.30 | -0.21 0.10 | 0.00 0.50 |
@@ 5.2 Chemistry in Aqu2
We compare our chemical abundances for Aqu2472 to Population III model yields from Heger & Woosley (2010, 2012 update). In Fig. 12, an excellent fit was found when combining the yields from two Pop III SN using StarFit212121StarFit: https://starfit.org/. These two stars included an 18.4 M⊙ SN with typical energy (2.4 x 1051 erg), combined with a 90 M⊙ SN with excess energy (10 x 1051 erg), both with mild mixing fractions. Ti was excluded from the fit in Fig. 12, as its NLTE corrections may be insufficient (e.g., see Mallinson et al., 2022, 2024); including Ti increases the value to 0.82. We note that our Ti abundance can be reproduced by including a third source, particularly with a significant r-process yield; however, the required additional model would also lead to an overproduction of Sr and Ba relative to our upper limit estimates, making such a scenario unlikely. For Aqu2776, none of the tested combinations of 2-3 Pop III SN yields provided a satisfactory fit (with ). Given its proximity to the system’s center, as shown in Fig. 8, and its overall higher metallicity, we attribute this to enrichment from more Pop III stars, and most likely pollution from super-AGB stars (discussed below in Section 5.2.1).
@@ 5.2.1 Element (anti-) correlations (K-Na)
Alternatively, the unusual globular cluste NGC 2419 is notable for its substantial scatter in K abundances, reaching unprecedented values of [K/Fe] dex (Cohen & Kirby, 2012; Mucciarelli et al., 2012); see Fig. 14. These high-K stars are found to be correlated with depleted Mg; i.e., a Mg-K anti-correlation. Ventura et al. (2012) proposed that hot-bottom burning in stars of masses around 6 M⊙ – at the edge between AGB and super-AGB (SAGB) regime – could reproduce the extreme potassium abundances in NGC 2419 if the standard cross section of Argon nuclei in reaction were increased by a factor of 100. In this model222222The SAGB model by Ventura et al. (2012) also predicts a modest increase in Na and a strong depletion in O, i.e., a weak Na-O anti-correlation similar to that observed in GCs. The SAGB scenario has been further supported by recent high-precision measurements of the potassium-destroying reaction (Fox et al., 2024), which is critically sensitive to temperature-density conditions (Dermigny & Iliadis, 2017)., the reduction of Mg and the production of K are maximized by also decreasing the mass loss rate by a factor of 4. By exploring a nuclear reaction network sensitive to variations in temperature, density, hydrogen abundance, reaction rates, and initial composition, Iliadis et al. (2016) identified that these abundance variations are limited to a narrow temperature-density range, achievable only by SAGB stars and classical novae.
@@ 6.2 Chemistry in Sgr2
We also examine the absolute abundances of Sgr2584 and Sgr2936232323Sgr2656 is not plotted due to having fewer abundance measurements, all of which are consistent with the other two targets. compared to solar values and our two standard stars, HD222925 and HD122563. In Fig. 19, the chemistry for the two standard stars in determined in two ways: (1) from the literature for HD222925 (Roederer et al., 2018) and HD122563 (Honda et al., 2006; Collet et al., 2018), and (2) derived using the same spectral lines as for the Sgr2 stars (details discussed further in Venn et al. 2025). The solar system’s s- and r-process abundance patterns from Simmerer et al. (2004) are scaled to match the Ba and Eu abundances in Sgr2584. The chemical abundances for both Sgr2 targets are similar, and generally fall between those of the two standard stars.
@@ 6.3 R-process Enrichments in Stellar Populations
As a reference globular cluster for comparison, M92 does not exhibit a clear gradient in Eu abundance with respect to radial distance or along the east-west and north-south directions (see Fig. 21). No gradient was seen in M15 either by (Cabrera Garcia et al., 2024, see their Fig. 13).
@@ 6.4 Comparison of Sgr2 with Tuc III
Tuc III is a particularly interesting system as it has been classified as a UFD based on its low average metallicity ([Fe/H] ; Li et al. 2018), a velocity and metallicity dispersion242424The observed metallicity gradient from CaT lines among core and tidal tail members has likewise been presented as evidence for the system’s origin as a UFD (Li et al., 2018). We reanalyze the evidence for a metallicity dispersion and metallicity gradients in the Appendix and do not find clear evidence for Tuc III to be classified as an UFD., and the absence of light-element anti-correlations amongst five of its brightest members (Hansen et al., 2017; Marshall et al., 2019). However, Marshall et al. (2019) note that adopting the photometric temperature for the metallicity outlier star from Hansen et al. (2017) would eliminate the statistically significant metallicity spread among its members. And, unfortunately, the velocity distribution for Tuc III is not representative of the system due to its disrupted state (only an upper limit of 1.5 km s-1 at the 95.5 % confidence level for core members was derived by Simon et al. 2017). Taking into account that low metallicity is also insufficient evidence for the presence of dark matter (i.e., an UFD classification), as globular clusters are typically found with [Fe/H] , and even lower ([Fe/H] have been found for M31 clusters Simpson 2018; Larsen et al. 2020, and the stellar stream C-19 appears to be a disrupted star cluster with [Fe/H] , see Martin et al. 2022; Yuan et al. 2022; Venn et al. 2025), then Tuc III remains a challenging system to classify.
@@ 7 Additional diagnostics for ambiguous systems
In addition to theoretical uncertainties, the observed velocity values can also be influenced by data quality and sample selection (e.g., compare the pink profile from L20 and the green profile from L21 in Fig. 17), as well as binary contamination. GC observations show that the binary fraction increases toward a cluster center (e.g., Sollima et al., 2007; Ji et al., 2015), and simulations reveal that as binaries segregate, "hard"252525”hard” means that their binding energy is much higher than the average kinetic energy of stars in a cluster binaries become even harder. Their semimajor axes shrink according to the Heggie 1975 - Hills 1975 law, making them more difficult to identify and potentially causing up to a 70% overestimation of the velocity dispersion in the cluster’s core (Aros et al., 2021). As a quantitative example, Wang et al. (2024) estimated the observed binary fraction for a Palomar 5-like globular cluster (which we note is close to Sgr2 on the vs. diagram; see Fig. 1), with a heliocentric distance of 20 kpc. They showed that most bright binaries with periods below days can be detected within 6 months by measuring line-of-sight velocities of km s-1. The undetected binaries ( km s-1) could still inflate the computed velocity dispersion by a factor of 1.5-2 compared to estimates based solely on single stars. For Sgr2, which is times farther away, the stars have measured radial velocity uncertainties 1 km s-1 (L21), constrained by observations of only 1 month (38 days). This implies its () is not resolved. This agrees with Baumgardt et al. (2022), who estimated a lower velocity dispersion for Sgr2262626The velocity dispersion profile for Sgr2 derived with N-body models of Baumgardt (2017) can be found at https://people.smp.uq.edu.au/HolgerBaumgardt/globular/fits/kin/sgrii_vel.pdf of km s-1, i.e., 3x lower than L21 (see Fig. 16), and attributed the higher observed () entirely to binary contamination.
To take these studies further, Baumgardt et al. (2022) have conducted a homogeneous and extensive study of mass segregation in GCs and UFD candidates using N-body simulations. As a measure of mass segregation, they used the ratio of the radii where the cumulative fraction of bright and faint stars reaches 0.5. A lower ratio (deviating further from unity) indicates stronger segregation and is expected in clusters with relaxation times much shorter than their age. To get the two subsets, members below the main sequence turn-off and above a certain MLow mass threshold were divided into two equally sized groups based on brightness. For Sgr2, their mass segregation parameter , which is in agreement with the expected value for a star cluster of Sgr2’s relaxation time of 8.37 Gyr, calculated following Spitzer (1987)272727We note that Wang et al. (2024) find that the presence of black holes can significantly shorten the two-component relaxation time..
@@
We support that the mass segregation parameter can provide an additional and valuable test for low surface brightness systems when trying to determine their origins (GC vs UFD), total mass, and thereby dark matter contents. The chemical patterns – particularly in heavy elements – and mass segregation of Sgr2 align with those of similar Milky Way GCs, supporting its classification as a star cluster. Such diagnostics are especially important when alternative approaches, like studying system stability within the Milky Way’s tidal field, are inapplicable or computationally expensive (e.g., Sgr2’s distant pericenter, as noted in Section 1, or the impact of the LMC on Tuc III).